# Rate-Distortion-Perception Tradeoff for Gaussian Vector Sources

### v1

@@ Appendix D Proof of Theorem 4
First, we show that the optimization problem in (101) is convex. The second derivative of the objective function (101a) with respect to $\gamma_{\ell}$ is $\frac{1}{2\gamma_{\ell}^{2}}$ which is positive. The second derivative of the function in the constraint (101e) with respect to $\hat{\lambda}_{\ell}$ is $\frac{1}{2\hat{\lambda}_{\ell}^{2}}$ which is again positive. It just remains to study the constraint (101d). The Hessian matrix of the function in this constraint is
@@
Since the $(D,P)$ is assumed to be strictly feasible, the Slater’s condition is satisfied. This implies that the solution to this problem is equal to that of the following dual optimization problem
$\max_{\nu_{1},\nu_{2},\eta_{\ell},\xi_{\ell}\geq 0}\;\;\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}$ $\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}+\nu_{1}\left(\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}})-D\right)$
$\quad+\nu_{2}\left(\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\right)+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell},$
@@
D-1 Case Where the Maximum for the Outer Optimization Occurs at $\nu_{1},\nu_{2}>0$
This is the case where both perception and distortion constraints are active. Let $\hat{\lambda}^{*}_{\ell}$ and $\gamma_{\ell}^{*}$ be the optimal solution to the inner minimization problem in (LABEL:Lagrange-dual-function) for the optimal $\nu_{1}$ and $\nu_{2}$. We first note that
@@ D-1 Case Where the Maximum for the Outer Optimization Occurs at $\nu_{1},\nu_{2}>0$
Next, we show that the following strict inequality holds:
@@
Suppose that the above strict inequality does not hold, i.e., $\gamma_{\ell}^{*}=\lambda_{\ell}$. We show that such $\gamma_{\ell}^{*}$ cannot be the optimal solution to the inner minimization problem.
The Lagrangian term in (LABEL:Lagrange-dual-function) depends on $\gamma_{\ell}$ and $\hat{\lambda}_{\ell}$ through the following function:
$G_{\ell}(\gamma_{\ell},\hat{\lambda}_{\ell})$ $=$ $\frac{1}{2}\log\frac{\lambda_{\ell}}{\gamma_{\ell}}+\nu_{1}\left(\lambda_{\ell}-2\sqrt{\hat{\lambda}_{\ell}(\lambda_{\ell}-\gamma_{\ell})}+\hat{\lambda}_{\ell}\right)+\frac{\nu_{2}}{2}\left(\frac{\hat{\lambda}_{\ell}}{\lambda_{\ell}}-1+\log\frac{\lambda_{\ell}}{\hat{\lambda}_{\ell}}\right)$ (138)
$\hskip 28.45274pt+\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\eta_{\ell}\hat{\lambda}_{\ell}.$
@@
$=$ $-\frac{\epsilon}{2\lambda_{\ell}}+2\nu_{1}\sqrt{\epsilon\hat{\lambda}^{*}_{{\ell}}}-\epsilon\xi_{\ell}+O(\epsilon^{2})$ (140)
$=$ $2\nu_{1}\sqrt{\epsilon\hat{\lambda}^{*}_{{\ell}}}+O(\epsilon)$ (141)
@@
$\frac{1}{2\gamma^{*}_{{\ell}}}-\nu_{1}\sqrt{\frac{{\hat{\lambda}_{{\ell}}^{*}}}{{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}}}-\xi_{\ell}$ $=$ $0,$ (142a)
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}{\hat{\lambda}^{*}_{{\ell}}}}+1\right)+\frac{1}{2}\nu_{2}\left(\frac{1}{\lambda_{{\ell}}}-\frac{1}{\hat{\lambda}^{*}_{{\ell}}}\right)-\eta_{\ell}$ $=$ $0,$ (142b)
$\xi_{\ell}(\gamma^{*}_{\ell}-\lambda_{\ell})$ $=$ $0,$ (142c)
$\eta_{\ell}\hat{\lambda}^{*}_{\ell}$ $=$ $0,$ (142d)
$\nu_{1}\left(\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}^{*}(\lambda_{{\ell}}-\gamma_{{\ell}}^{*})}+\hat{\lambda}_{{\ell}}^{*}\right)-D\right)$ $=$ $0,$ (142e)
$\nu_{2}\left(\sum_{\ell=1}^{L}\frac{1}{2}\left(\frac{\hat{\lambda}_{{\ell}}^{*}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}^{*}}\right)-P\right)$ $=$ $0,$ (142f)
@@
Due to the strict inequalities (137) and (136), we have that $\xi_{\ell}=0$ and $\eta_{\ell}=0$. Then, from condition (142a), we can write $\hat{\lambda}^{*}_{\ell}$ as follows
@@
Note that as $\gamma^{*}_{\ell}$ varies from 00 to $\lambda_{\ell}$, the left-hand side of (144) decreases monotonically from $\nu_{1}$ to $(1-2\nu_{1}\lambda_{\ell})\nu_{1}$ while the right-hand side of (144) increases monotonically from $-\frac{\nu_{2}}{2\lambda_{\ell}}$ to +$\infty$+ ∞ So, this equation has a unique solution in the interval $(0,\lambda_{\ell})$. The equation (144) is quadratic, so it can solved analytically. The solution gives (IV-C) and (37).
D-2 Case Where the Maximum for the Outer Optimization Occurs at $\nu_{1}>0,\nu_{2}=0$
This is the case where the distortion metric is active but the perception metric is inactive. Clearly, this reduces to the traditional rate-distortion function.
D-3 Case Where the Maximum for the Outer Optimization Occurs at $\nu_{1}=0$
This is the case where the distortion metric is inactive, so the inner minimization problem in (LABEL:Lagrange-dual-function) decouples into two independent minimizations, one for $\gamma_{\ell}$ and the other one for $\hat{\lambda}_{\ell}$, i.e.,
$\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}+\nu_{2}\left(\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\right)+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}$
$\quad=\min_{\{\gamma_{\ell}\}_{\ell=1}^{L}}\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})$
$\qquad+\min_{\{\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\nu_{2}\left(\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\right)-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}.$ (145)
@@ Appendix E Proof of Theorem 5
The lower bounding steps for the perception metric are as follows
$W_{2}^{2}(P_{X},P_{\hat{X}^{*}_{G}})$ $=$ $\text{tr}(\Sigma_{X}+\Sigma_{\hat{X}^{*}_{G}}-2(\Sigma_{X}^{\frac{1}{2}}\Sigma_{\hat{X}^{*}_{G}}\Sigma_{X}^{\frac{1}{2}})^{\frac{1}{2}})$ (153)
$=$ $\text{tr}(\Theta\Sigma_{X}\Theta^{T}+\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}-2\Theta(\Sigma_{X}^{\frac{1}{2}}\Sigma_{\hat{X}^{*}_{G}}\Sigma_{X}^{\frac{1}{2}})^{\frac{1}{2}}\Theta^{T})$ (154)
$=$ $\text{tr}(\Theta\Sigma_{X}\Theta^{T}+\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}-2(\Theta\Sigma_{X}^{\frac{1}{2}}\Sigma_{\hat{X}^{*}_{G}}\Sigma_{X}^{\frac{1}{2}}\Theta^{T})^{\frac{1}{2}})$ (155)
$=$ $\text{tr}(\Theta\Sigma_{X}\Theta^{T}+\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}-2(\Theta\Sigma_{X}^{\frac{1}{2}}\Theta^{T}\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}\Theta\Sigma_{X}^{\frac{1}{2}}\Theta^{T})^{\frac{1}{2}})$ (156)
$=$ $\text{tr}(\Theta\Sigma_{X}\Theta^{T}+\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}-2((\Theta\Sigma_{X}\Theta^{T})^{\frac{1}{2}}\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T}(\Theta\Sigma_{X}\Theta^{T})^{\frac{1}{2}})^{\frac{1}{2}})$ (157)
$=$ $W_{2}^{2}(P_{\Theta X},P_{\Theta\hat{X}^{*}_{G}})$ (158)
@@
$=$ $\sum_{\ell=1}^{L}(\sqrt{\mathbbm{E}[(Z_{\ell})^{2}]}-\sqrt{\mathbbm{E}[(\hat{Z}^{*}_{G,\ell})^{2}]})^{2}$ (161)
@@
  • (154) follows because the trace is invariant under unitary transformations;
  • (155) and (157) follow because for a given matrix $A$, $(\Theta A\Theta^{T})^{\frac{1}{2}}=\Theta A^{\frac{1}{2}}\Theta^{T}$ since $\Theta$ is a unitary matrix;
  • (156) follows because $\Theta^{T}\Theta=I$;
@@
$\hskip 113.81102pt0\leq\hat{\lambda}_{\ell},$ (163c)
$\hskip 113.81102pt\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}}\right)\leq D,$ (163d)
$\hskip 113.81102pt\sum_{\ell=1}^{L}\left(\sqrt{\lambda_{{\ell}}}-\sqrt{\hat{\lambda}_{{\ell}}}\right)^{2}\leq P.$ (163e)
@@ Appendix F Proof of Theorem 6
$\max_{\nu_{1},\nu_{2},\eta_{\ell},\xi_{\ell}\geq 0}\;\;\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}$ $\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}+\nu_{1}\left(\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}})-D\right)$ (165)
$+\nu_{2}\left(\sum_{\ell=1}^{L}\left(\sqrt{\lambda_{{\ell}}}-\sqrt{\hat{\lambda}_{{\ell}}}\right)^{2}-P\right)+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}.$
@@
$G^{\prime}_{\ell}(\gamma_{\ell},\hat{\lambda}_{\ell})$ $=$ $\frac{1}{2}\log\frac{\lambda_{\ell}}{\gamma_{\ell}}+\nu_{1}\left(\lambda_{\ell}-2\sqrt{\hat{\lambda}_{\ell}(\lambda_{\ell}-\gamma_{\ell})}+\hat{\lambda}_{\ell}\right)+\nu_{2}\left(\sqrt{\lambda_{\ell}}-\sqrt{\hat{\lambda}_{\ell}}\right)^{2}$ (166)
$\hskip 28.45274pt+\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\eta_{\ell}\hat{\lambda}_{\ell}.$
@@
$G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{*}_{\ell})-G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{\prime}_{\ell})$ $=$ $\nu_{1}(2\sqrt{\epsilon(\lambda_{\ell}-\gamma_{\ell}^{*})}-\epsilon)+\nu_{2}(2\sqrt{\lambda_{\ell}\epsilon}-\epsilon)+\eta_{\ell}\epsilon$ (167)
@@
Thus, if $\nu_{2}>0$, for sufficiently small $\epsilon>0$, we can strictly decrease $G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{*}_{\ell})$, while satisfying the implicit constraints. This contradicts with the assumption that $\hat{\lambda}^{*}_{\ell}=0$ is the optimal solution to the inner minimization problem. This proves (164). Given the strict inequality in (164), similar to the KL-divergence metric, we can show that
@@
$\frac{1}{2\gamma^{*}_{{\ell}}}-\nu_{1}\sqrt{\frac{{\hat{\lambda}_{{\ell}}^{*}}}{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}}$ $=$ $0,$ (170a)
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}{\hat{\lambda}^{*}_{{\ell}}}}+1\right)+\nu_{2}\left(1-\sqrt{\frac{\lambda_{\ell}}{\hat{\lambda}^{*}_{\ell}}}\right)$ $=$ $0,$ (170b)
$\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}^{*}(\lambda_{{\ell}}-\gamma_{{\ell}}^{*})}+\hat{\lambda}_{{\ell}}^{*})$ $=$ $D,$ (170c)
@@
The derivation of the optimal solution can now be shown as follows. Define
@@
$\hat{\lambda}^{*}_{\ell}$ $=$ $\frac{\lambda_{\ell}}{\left(1+\frac{(1-\theta_{\ell})\nu_{1}}{\nu_{2}}\right)^{2}},$ (172)
@@
Note that the function $\frac{\theta_{\ell}}{1+\frac{(1-\theta_{\ell})\nu_{1}}{\nu_{2}}}$ is an increasing function in $\theta_{\ell}$. Also, the function $\sqrt{1-\frac{\theta_{\ell}}{2\nu_{1}\lambda_{\ell}}}$ as defined in $\theta_{\ell}\in[0,2\nu_{1}\lambda_{\ell}]$ is a decreasing function in $\theta_{\ell}$. So, the solution to the above equation is unique.

### v2

@@ Appendix D Proof of Theorem 4
First, we show that the optimization problem in (105) is convex. The second derivative of the objective function (105a) with respect to $\gamma_{\ell}$ is $\frac{1}{2\gamma_{\ell}^{2}}$ which is positive. The second derivative of the function in the constraint (105e) with respect to $\hat{\lambda}_{\ell}$ is $\frac{1}{2\hat{\lambda}_{\ell}^{2}}$ which is again positive.
It remains to study the constraint (105d). The Hessian matrix of the function in this constraint is
@@
Since $(D,P)$ is assumed to be strictly feasible, the Slater’s condition is satisfied. This implies that the solution to this problem is equal to that of the following dual optimization problem
$\max_{\nu_{1},\nu_{2},\eta_{\ell},\xi_{\ell}\geq 0}\;\;\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}$
$\hskip 28.45274pt+\nu_{1}\left(\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}})-D\right)$
$\hskip 28.45274pt+\nu_{2}\left(\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\right)$
$\hskip 28.45274pt+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell},$ (145)
@@
D-A Maximum of Outer Optimization Occurs at $\nu_{1},\nu_{2}>0$
This is the case where both perception and distortion constraints are active. Let $\hat{\lambda}^{*}_{\ell}$ and $\gamma_{\ell}^{*}$ be the optimal solution to the inner minimization problem in (145) for the optimal $\nu_{1}$ and $\nu_{2}$.
We first note that
@@ D-A Maximum of Outer Optimization Occurs at $\nu_{1},\nu_{2}>0$
Next, we prove that the following strict inequality holds:
@@
We proceed with proof by contradiction. Suppose that the above strict inequality does not hold, i.e., $\gamma_{\ell}^{*}=\lambda_{\ell}$. We show that such $\gamma_{\ell}^{*}$ cannot be the optimal solution to the inner minimization problem.
$G_{\ell}(\gamma_{\ell},\hat{\lambda}_{\ell})$ $=$ $\frac{1}{2}\log\frac{\lambda_{\ell}}{\gamma_{\ell}}+\nu_{1}\left(\lambda_{\ell}-2\sqrt{\hat{\lambda}_{\ell}(\lambda_{\ell}-\gamma_{\ell})}+\hat{\lambda}_{\ell}\right)$ (148)
$\hskip 14.22636pt+\frac{\nu_{2}}{2}\left(\frac{\hat{\lambda}_{\ell}}{\lambda_{\ell}}-1+\log\frac{\lambda_{\ell}}{\hat{\lambda}_{\ell}}\right)$
$\hskip 14.22636pt+\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\eta_{\ell}\hat{\lambda}_{\ell}.$
@@
$=$ $-\frac{\epsilon}{2\lambda_{\ell}}+2\nu_{1}\sqrt{\epsilon\hat{\lambda}^{*}_{{\ell}}}-\epsilon\xi_{\ell}+O(\epsilon^{2})$
$=$ $2\nu_{1}\sqrt{\epsilon\hat{\lambda}^{*}_{{\ell}}}+O(\epsilon),$
@@
[…]
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}{\hat{\lambda}^{*}_{{\ell}}}}+1\right)+\frac{1}{2}\nu_{2}\left(\frac{1}{\lambda_{{\ell}}}-\frac{1}{\hat{\lambda}^{*}_{{\ell}}}\right)-\eta_{\ell}$ $=$ 00 0\;\;\; (152b)
$\xi_{\ell}(\gamma^{*}_{\ell}-\lambda_{\ell})$ $=$ 00 0\;\;\; (152c)
$\eta_{\ell}\hat{\lambda}^{*}_{\ell}$ $=$ 00 0\;\;\; (152d)
$\nu_{1}\left(\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}^{*}(\lambda_{{\ell}}-\gamma_{{\ell}}^{*})}+\hat{\lambda}_{{\ell}}^{*}\right)-D\right)$ $=$ 00 0\;\;\; (152e)
$\nu_{2}\left(\sum_{\ell=1}^{L}\frac{1}{2}\left(\frac{\hat{\lambda}_{{\ell}}^{*}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}^{*}}\right)-P\right)$ $=$ 00 0\;\;\; (152f)
@@
Due to the strict inequalities (147) and (146), we have that $\xi_{\ell}=0$ and $\eta_{\ell}=0$. Then, from condition (152a), we can write $\hat{\lambda}^{*}_{\ell}$ as follows:
@@
Note that as $\gamma^{*}_{\ell}$ varies from 00 to $\lambda_{\ell}$, the left-hand side of (154) decreases monotonically from $\nu_{1}$ to $(1-2\nu_{1}\lambda_{\ell})\nu_{1}$ while the right-hand side of (154) increases monotonically from $-\frac{\nu_{2}}{2\lambda_{\ell}}$ to +$\infty$+ ∞. So, this equation has a unique solution in the interval $(0,\lambda_{\ell})$. The equation (154) is quadratic, so it can be solved analytically. The solution gives (IV-C) and (39).
D-B Maximum of Outer Optimization Occurs at $\nu_{1}>0,\nu_{2}=0$
This is the where the distortion metric is active, but the perception metric is inactive. Clearly, this reduces to the traditional rate-distortion function.
D-C Maximum of Outer Optimization Occurs at $\nu_{1}=0$
This is the case where the distortion metric is inactive, so the inner minimization problem in (145) decouples into two independent minimizations, one for $\gamma_{\ell}$ and the other one for $\hat{\lambda}_{\ell}$, as shown below
$\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\left\{\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}\right.$
$\qquad\qquad\qquad+\nu_{2}\left(\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\right)$
$\qquad\qquad\qquad\left.+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}\right\}$
$=\min_{\{\gamma_{\ell}\}_{\ell=1}^{L}}\left\{\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})\right\}$
$\qquad+\min_{\{\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\left\{\nu_{2}\Bigg{(}\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)-P\Bigg{)}\right.$
$\qquad\qquad\qquad\qquad\left.-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}\right\}.$ (155)
@@ Appendix E Proof of Theorem 5
The lower bounding steps for the perception metric are as follows:
$W_{2}^{2}(P_{X},P_{\hat{X}^{*}_{G}})$ $=$ $W_{2}^{2}(P_{\Theta X},P_{\Theta\hat{X}^{*}_{G}})$ (163)
@@
$=$ $\sum_{\ell=1}^{L}\left(\sqrt{\mathbbm{E}[(Z_{\ell})^{2}]}-\sqrt{\mathbbm{E}[(\hat{Z}^{*}_{G,\ell})^{2}]}\right)^{2}\quad\;\;$ (166)
@@
  • (163) follows because $\Theta$ is a unitary matrix and Wasserstein-2 distance is invariant under unitary transformations;
@@
$R(D,P)$ (168e)
$=$ $\min_{\{\hat{\lambda}_{\ell},\gamma_{{\ell}}\}_{\ell=1}^{L}}\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}$
@@
$\hskip 39.83368pt0\leq\hat{\lambda}_{\ell}$
$\hskip 39.83368pt\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}}\right)\leq D\qquad\;\;$
$\hskip 39.83368pt\sum_{\ell=1}^{L}\left(\sqrt{\lambda_{{\ell}}}-\sqrt{\hat{\lambda}_{{\ell}}}\right)^{2}\leq P.$
@@ Appendix F Proof of Theorem 6
$\max_{\nu_{1},\nu_{2},\eta_{\ell},\xi_{\ell}\geq 0}\;\;\min_{\{\gamma_{\ell},\hat{\lambda}_{\ell}\}_{\ell=1}^{L}}\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}$
$\hskip 28.45274pt+\nu_{1}\left(\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}})-D\right)$
$\hskip 28.45274pt+\nu_{2}\left(\sum_{\ell=1}^{L}\left(\sqrt{\lambda_{{\ell}}}-\sqrt{\hat{\lambda}_{{\ell}}}\right)^{2}-P\right)$
$\hskip 28.45274pt+\sum_{\ell=1}^{L}\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\sum_{\ell=1}^{L}\eta_{\ell}\hat{\lambda}_{\ell}.$ (170)
@@
$G^{\prime}_{\ell}(\gamma_{\ell},\hat{\lambda}_{\ell})$ $=$ $\frac{1}{2}\log\frac{\lambda_{\ell}}{\gamma_{\ell}}+\nu_{1}\left(\lambda_{\ell}-2\sqrt{\hat{\lambda}_{\ell}(\lambda_{\ell}-\gamma_{\ell})}+\hat{\lambda}_{\ell}\right)$
$\hskip 2.84544pt+\nu_{2}\left(\sqrt{\lambda_{\ell}}-\sqrt{\hat{\lambda}_{\ell}}\right)^{2}+\xi_{\ell}(\gamma_{\ell}-\lambda_{\ell})-\eta_{\ell}\hat{\lambda}_{\ell}.$
@@
$G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{*}_{\ell})-G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{\prime}_{\ell})$ (172)
$=$ $\nu_{1}(2\sqrt{\epsilon(\lambda_{\ell}-\gamma_{\ell}^{*})}-\epsilon)+\nu_{2}(2\sqrt{\lambda_{\ell}\epsilon}-\epsilon)+\eta_{\ell}\epsilon\quad$
@@
Thus, if $\nu_{2}>0$, for sufficiently small $\epsilon>0$, we can strictly decrease $G^{\prime}_{\ell}(\gamma^{*}_{\ell},\hat{\lambda}^{*}_{\ell})$, while satisfying the implicit constraints. This contradicts with the assumption that $\hat{\lambda}^{*}_{\ell}=0$ is the optimal solution to the inner minimization problem. This proves (169).
Given the strict inequality in (169), similar to the KL-divergence metric, we can show that
@@
$\frac{1}{2\gamma^{*}_{{\ell}}}-\nu_{1}\sqrt{\frac{{\hat{\lambda}_{{\ell}}^{*}}}{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}}$ $=$ 00 0$\quad$ (175a)
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}}{\hat{\lambda}^{*}_{{\ell}}}}+1\right)+\nu_{2}\left(1-\sqrt{\frac{\lambda_{\ell}}{\hat{\lambda}^{*}_{\ell}}}\right)$ $=$ 00 0$\quad$ (175b)
$\sum_{\ell=1}^{L}(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}^{*}(\lambda_{{\ell}}-\gamma_{{\ell}}^{*})}+\hat{\lambda}_{{\ell}}^{*})$ $=$ $D\quad$ (175c)
@@
The derivation of the optimal solution can now be shown as follows.
Define
@@
$\hat{\lambda}^{*}_{\ell}$ $=$ $\frac{\lambda_{\ell}}{\left(1+\frac{(1-\theta_{\ell})\nu_{1}}{\nu_{2}}\right)^{2}}.$ (177)
@@
Note that the function $\frac{\theta_{\ell}}{1+\frac{(1-\theta_{\ell})\nu_{1}}{\nu_{2}}}$ is an increasing function in $\theta_{\ell}$. Also, the function $\sqrt{1-\frac{\theta_{\ell}}{2\nu_{1}\lambda_{\ell}}}$ in the interval $\theta_{\ell}\in[0,2\nu_{1}\lambda_{\ell}]$ is a decreasing function in $\theta_{\ell}$. So, the solution to the above equation is unique.